# CHARTOM: A Visual Theory-of-Mind Benchmark for Multimodal Large Language Models

### v2

@@ 4.2 Certain Manipulations are More Misleading Than Others
We compute the log odds ratio of making mistakes by humans and LLMs while answering FACT questions on non-manipulated vs. manipulation conditions, and show the results in Figure 5. The log odds-ratio is computed as follows:
@@
where $\hat{p}_{m},\hat{p}_{n}$ are the maximum likelihood estimates of making mistakes on manipulated and non-manipulated charts, respectively, by a population of agents. A higher odds-ratio suggests that manipulation is more misleading. We report the log odds ratio for each group and manipulation type.
To estimate $\hat{p}_{m}$ and $\hat{p}_{n}$, we fit separate mixed-effects logistic regression models for human and LLM responses. For LLMs, the model included fixed effects for group id and manipulation type, along with random intercepts for individual LLMs and question id within each group id. Similarly, for human responses, the model included fixed effects for group id and manipulation type, with random intercepts for human id and question id within each group. Estimated marginal means from these models were used to obtain $\hat{p}_{m}$ and $\hat{p}_{n}$ values which were then used for computing log-odds ratio.
@@
    Manipulations that effectively mislead humans tend to also deceive LLMs: There is a moderate to strong correlation (Kendall’s Tau score of 0.60) between the rankings of 14 groups based on their odds ratio in humans vs LLMs.
    • Top manipulation groups like line+manipulated x-axis, line+inverted y-axis, scatter + swapped axes, scatter+logarithmic y-axis are significantly misleading to both.
    • Bottom manipulations like bar+truncated y-axis, bar+compressed y-axis do not mislead humans. However, some of them like bar+pictorial bar and line+compressed y-axis, pie+3D still mislead LLMs, suggesting different modes of failure.
@@
    The two groups have significantly different scales of log odds ratio: The log odds ratio of LLMs on top manipulations in Figure 5 is on the order of $10^{20}$, compared to that of humans which is roughly on the order of $10^{5}$. This shows that when LLMs are misled, they are misled to a significant extent further. This again suggests different modes of failure.
4.3 LLMs Failed to Estimate Chart Misleadingness (HMI)
We asked each LLM the MIND question (what fraction of humans will be misled) on each chart. Table 2 reports the mean absolute error (MAE) between the LLM-predicted and ground truth HMI scores.
LLM Claude Gemini GPT-4o Llama LLaVA Qwen
MAE 0.40 $\pm$ 0.02 0.37 $\pm$ 0.02 0.35 $\pm$ 0.02 0.46 $\pm$ 0.02 0.32 $\pm$ 0.02 0.29 $\pm$ 0.02
Table 2: MAE of LLMs on MIND questions
The large MAE values suggest that the LLMs lacked certain visual theory-of-mind: they cannot reliably predict the extent to which humans are misled by various chart manipulations.
Figure 6: MIND Performance: LLM predicted HMI lacks the distinction shown in human ground truth with respect to high/low groups.
LLMs lack discriminative power for predicting HMI scores for high and low true HMI charts.
We further grouped all the charts in our benchmark dataset into high-HMI charts (charts with HMI0.5absent0.5$\geq$ 0.5≥ 0.5) and low-HMI charts (those with HMI<0.5). We computed the average predicted HMI by each LLM on both groups and plotted the results in Figure 6. For reference, we show the ground truth HMI on the left. We make the following observations:
@@ 4.3 LLMs Failed to Estimate Chart Misleadingness (HMI)
    LLMs show limited discriminative power on MIND questions: While all LLMs predict high-HMI charts to be more misleading than low-HMI ones on average, the difference is much smaller compared to the gap seen in human ground truth on the left of Figure 6. This indicates that LLMs lack a good discriminative power on MIND questions.
@@
    LLMs under/over estimate HMI of high and low HMI charts: As a corollary, we observe that LLMs consistently underestimate HMI for high misleading charts(orange bars of LLMs are consistently lower than orange ground truth HMI bar of human on the left) and overestimate the HMI for low misleading charts(blue bars follow the opposite trend instead).
@@ 5 Conclusion
We presented a new theory-of-mind benchmark called CHARTOM for LLMs based on visual chart comprehension. Our benchmark contained two types of questions: FACT questions based on factual information, and, MIND questions based on human theory of mind. We obtained the ground truth for MIND questions from our human experiments. Finally, we evaluated the performance of leading LLMs on both FACT and MIND questions and reported our findings. In summary, we observed that these LLMs perform poorly on both FACT and MIND questions indicating substantial room for improvement. We hope our benchmark will contribute to the development of future LLMs.

### v3

@@ 5.2 Certain Manipulations are More Misleading Than Others
Different manipulations have different “degree” by which they mislead a population of humans or LLMs. We characterize this “degree” by log odds ratio of making mistakes by a population on manipulated charts when compared to non-manipulated ones.
Log odds ratio
: The log odds ratio characterizes the degree of misleadingness of a manipulation is computed as follows:
@@
where, $\hat{p}_{m},\hat{p}_{n}$ are the maximum likelihood estimates of making mistakes on manipulated and non-manipulated charts, respectively, by a population of agents. A higher odds ratio suggests that manipulation is more misleading.
We report the log odds ratio for each manipulation group by both humans and LLMs in Figure 4. To estimate $\hat{p}_{m}$ and $\hat{p}_{n}$, we fit separate mixed-effects logistic regression models on human and LLM responses. For both human and LLM populations, the model considers group id and manipulation type as fixed effects, and agent id and question id within each group id as random effects. The estimated marginal means from these models were used to obtain $\hat{p}_{m}$ and $\hat{p}_{n}$ values which were then used for computing the log-odds ratio in equation 1.
@@
    Manipulations that effectively mislead humans tend to also deceive LLMs: In Figure 4, we show different manipulations in decreasing order of their “degree" of misleadingness to humans along with corresponding log odds ratio for both humans and LLMs. We observe a moderate to strong correlation (Kendall’s Tau score of 0.60) between the rankings of 14 manipulations between humans and LLMs. Furthermore, we note that,
    • Top manipulation like line+manipulated x-axis, line+inverted y-axis, scatter + swapped axes, scatter+logarithmic y-axis are significantly misleading to both humans and LLMs.
    • Bottom manipulations like bar+truncated y-axis, bar+compressed y-axis do not mislead humans. However, some of them like bar+pictorial bar and line+compressed y-axis, pie+3D still mislead LLMs though less strongly.
@@
    The two groups have significantly different scales of log odds ratio: The log odds ratio of LLMs on top manipulations in Figure 4 is on the order of $10^{20}$, compared to that of humans which is roughly on the order of $10^{5}$. This shows that when LLMs are misled, they are misled much more significantly. This suggests different modes of failure.
5.3 LLMs Failed to Estimate Chart Misleadingness to Humans (HMI)
We asked each LLM the MIND question (what fraction of humans will be misled) on each chart. Table 2 reports the mean and standard deviation of the absolute error between the LLM-predicted and ground truth HMI scores.
gemini-2.5 qwen-2 qwen-2.5 llava gpt-o1 gpt-4o gemini-1.5 claude-3.7 claude-3.5 llama-3 MAE 0.24 0.29 0.30 0.32 0.32 0.35 0.37 0.37 0.40 0.46 $\hskip 7.11317pt\sigma$ 0.16 0.18 0.22 0.18 0.20 0.22 0.24 0.23 0.22 0.27
Table 2: MAE of predicting HMI(MIND answer) by LLMs.
We also show the descriptive performance of each LLM on the MIND question through scatter plots in Figure 5. The x-axis represents the ground truth HMI of all charts sorted in increasing order. A perfect prediction should align the dots along $y=x$ lines. We fit a linear curve on the prediction made by LLMs and show it by green line. We also show Pearson’s degree of correlation between actual HMI and predicted HMI on lower end corner for each LLM.
We make the following observations based on these results:
@@ 5.3 LLMs Failed to Estimate Chart Misleadingness to Humans (HMI)
    LLMs lack visual theory of mind on charts: This is evident from the large MAE values for each LLM which suggests that they cannot reliably predict the extent to which humans are misled by various chart manipulations.
@@
    LLMs show limited discriminative power between high and how HMI charts: Most LLMs except gemini-2.5 show almost a flat trend(as shown by the green line in Figure 5) in predicting HMI. This indicates that they can’t effectively differentiate between low and high HMI charts. However, gemini-2.5 shows some linear trend with a Person’s correlation of 0.6 suggesting that it can predict misleading aspects in charts much better.
    As a corollary, we remark that LLMs over or under-estimate HMI of ground truth low or high HMI charts respectively. This trend is also evident from the per manipulation arrow plot presented in Figure 7 in the appendix.
Figure 5: LLMs prediction vs ground truth on MIND questions.
6 Impacts & Limitations
Our work takes an important step towards understanding how LLMs understand and interpret misleading charts which is an area of great societal value. By releasing this novel benchmark, we lay the foundation for developing future LLMs that can not only recognize misleading charts but can also serve as useful assistants to warn and educate human users against misinformation to curb them.
A primary limitation of our benchmark is the relatively small size of our data, which is due in part to the high cost of conducting human experiments in a controlled setting. Additionally, our dataset mainly focuses on real-world charts that can be programmatically generated. As a result, it does not capture the full diversity of chart formats that may appear in the wild, including those created by non-standard means.
@@ 7 Conclusion
We presented a new theory-of-mind benchmark called CHARTOM for LLMs based on comprehension of misleading charts. Our benchmark contained two types of questions, one that tested understanding of LLMs on factual content present in charts, and, the other tested understanding of LLM on how much and why a chart is misleading to the human mind.
We evaluated the performance of leading LLMs on both FACT and MIND questions and observed that current LLMs not only possess a poor understanding of the degree of misleadingness of charts to humans but also fail to answer factual questions on manipulated charts thereby suggesting a big scope of improvement in future LLMs to understand and help humans with misleading charts.
@@ C Logistic Regression Models for Estimating $\hat{p}$
8 Details on Human Experiment
To compute the Human Misleadingness Index (HMI), we collected data from 78 undergraduate participants at a Midwestern U.S. university. Participants completed a within-subject experiment in which they viewed 112 visualizations—56 with manipulated visualizations and 56 corresponding non-manipulated counterparts—representing 14 types of manipulations (e.g., truncated y-axis, inverted y-axis, dual axes). For each visualization, participants answered a question requiring data interpretation, such as comparing values or identifying trends. The study began with informed consent and an instructional video simulating real-world scenarios (e.g., scrolling through graphs on social media), after which participants completed the visualization interpretation tasks at their own pace. Sanity-check questions were included to filter inattentive responses, and 10 participants were excluded based on this criterion. The final dataset from 68 participants was used to compute accuracy-based scores, quantifying the deceptive impact of each manipulation type as perceived by human viewers.
A screenshot of chart with questions presented to human subjects are given below(chart id is provided here just for reference):
Each participant received a \$20 Amazon gift card as compensation for their time and effort.
@@ 8 Details on Human Experiment
We present some selected chart and FACT question pairs from each manipulation group below.